OpenStudy (anonymous):

is f(x)=2x + 5 neither even nor odd?

OpenStudy (anonymous):

ODD: f(-x)=f(x) even: f(-x)=f(x)

OpenStudy (anonymous):

if they are not equal they are neither

OpenStudy (anonymous):

substitute -x in the function

OpenStudy (anonymous):

if you substitute x the answer is neither the exact opposite nor the original answer, I was looking for verification that it's neither even nor odd

OpenStudy (anonymous):

*-x

OpenStudy (anonymous):

oh a typo sorry ODD= -f(x)=f(-x) -------------- Substitute -x in the equation.. Multiply - to the equation.

OpenStudy (anonymous):

a polynomial is only odd if the degrees of ALL the terms are odd the degree of \(x\) is \(1\) which is odd, but the degree of the constant \(5\) is \(0\) which is even

OpenStudy (anonymous):

THANK YOU I was trying to remember that rule, but instead of considering DEGREES I was considering EXPONENTS that's VERY VERY helpful of you!

OpenStudy (anonymous):

yw